OpenStudy (anonymous):

Find the values of x so that the series below converges. (((x+8)^(n))/(2^n)) please help find the interval. goes to INF and n=0

OpenStudy (thomas5267):

Is this the equation? \[ a_n=\frac{(x+8)^n}{2^n} \]

OpenStudy (anonymous):

yes

OpenStudy (thomas5267):

What is the required limit?

OpenStudy (anonymous):

I'm looking to find the values of x so that the series below converges. (the interval it does

OpenStudy (anonymous):

limit is n=0

OpenStudy (anonymous):

to INF

OpenStudy (anonymous):

does that makes sense can u help?

OpenStudy (thomas5267):

So the question is find x such that: \[ \lim_{n\to0}\frac{(x+8)^n}{2^n}=\infty \]

OpenStudy (anonymous):

no it has the simga logo in front of the fraction with a INF aboe th sigma and a n=0 below

OpenStudy (anonymous):

\[\sum_{n=0}^{INF}\]

OpenStudy (thomas5267):

Find such x such that: \[ \sum_{n=0}^\infty\frac{(x+8)^n}{2^n}=\infty \]